OpenStudy (anonymous):

loga X + loga (x-2)=loga(x+4) Help please! step by step.

OpenStudy (anonymous):

Please pot the full question

OpenStudy (anonymous):

well its just saying to solve the equation.

OpenStudy (anonymous):

Log X + Log(x-2) = Log(x+4) Rule: Log(a.b)= Log a + Log b Therefore: Log x.(x-2) = Log(x+4) = Log x^2-2x = Log(x+4) Rule: Log(a)^x= xLog a Log (a/b) = Log a-Log b Therefore: 2Log x/ Log 2x = Log(x+4) = x^2/2x = (x+4) x^2= (x+4)2x x^2 = 2x^2*8x x^2 - 2x^2* = 8x -(x^2) = 8x x = -8 Please confirm the answer by substituting x